You have a portfolio that is invested 18 percent in Stock R, 40 percent in Stock S, and the remainder in Stock T. The beta of Stock R is .63, and the beta of Stock S is 1.18. The beta of your portfolio is 1.29. What is the beta of the Stock T?
Answer #1

Investment in T=100-(18+40)=42%

Portfolio beta=Respective beta*Respective investment weight

1.29=(0.18*0.63)+(0.4*1.18)+(0.42*beta of the Stock T)

1.29=.5854+(0.42*beta of the Stock T)

beta of the Stock T=(1.29-0.5854)/0.42

=1.68(Approx).